a rectangular dog kennel made using the wall of a barn and 24m of wire (to form the other three sides of the kennel). if the area of the kennel is 70m squared, find the dimensions
Answer by checkley71(8403)

X*Y=70
2Y+X=24 THUS X=24-2Y SUBSTITUTING 24-2Y FOR X IN THE FIRST FORMULA WE GET
(24-2Y)Y=70 OR 24Y-2Y~2-70=0 OR 2Y~2-24Y+70=O OR Y~2-12Y+35 OR
(Y-7)(Y-5)=0 OR Y=7 & Y=5 SUBSTITUTING 7 FOR Y IN EQUATION 1
7*X=70 OR X=70/10 OR X=7 THUS 2*7+10=24 THEN 24=24
5*X=70 OR X=70/5 OR X=14 THUS 2*5+14=24 THEN 24=24
THERE ARE TWO SETS OF SOLUTIONS X=7,Y=10 & X=5,Y=14